evaluate the following integral
∫ 1 – sin 2x / x + cos^2 (x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{1-sin2x}{x+cos^{2}x}dx
x+cos^{2}x = t
(1+2cosx(-sinx))dx = dt
(1-sin2x)dx = dt
I = \int \frac{dt}{t}
I = log(t)+c
I = log(1-sin2x)+c